if (1.200x10^2)g of methanol is burnt, calculate the energy change. Express tour answers in kilojoules, with the correct sign for a combustion reaction
Answer #1

​​​​​​As When one mole of methanol is burnt molar heat of combustion is 726 KJ/mole. Therefore ∆H reaction = -726 kJ/ mole.

As one mole of methanol = 32 gram( molecular weight of methanol.

32 g of methanol burnt release energy i.e.∆H = -726KJ/mole

1 g of metanolm burnt release energy = -726KJ mol-1/32g = -22.6875 KJ/mol-1g-1

1.200*10^2 g of methanol burnt release energy= (1.200*10^2)(-22.6875) = -2722.5KJ/mol